# A Survey on Large Language Models with Multilingualism: Recent Advances and New Frontiers

### v1

@@ Abstract
The rapid development of Large Language Models (LLMs) demonstrates remarkable multilingual capabilities in natural language processing, attracting global attention in both academia and industry. To mitigate potential discrimination and enhance the overall usability and accessibility for diverse language user groups, it is important for the development of language-fair technology. Despite the breakthroughs of LLMs, the investigation into the multilingual scenario remains insufficient, where a comprehensive survey to summarize recent approaches, developments, limitations, and potential solutions is desirable. To this end, we provide a survey with multiple perspectives on the utilization of LLMs in the multilingual scenario. We first rethink the transitions between previous and current research on pre-trained language models. Then we introduce several perspectives on the multilingualism of LLMs, including training and inference methods, model security, multi-domain with language culture, and usage of datasets. We also discuss the major challenges that arise in these aspects, along with possible solutions. Besides, we highlight future research directions that aim at further enhancing LLMs with multilingualism. The survey aims to help the research community address multilingual problems and provide a comprehensive understanding of the core concepts, key techniques, and latest developments in multilingual natural language processing based on LLMs.
@@ 1 Introduction
With the rapid development of artificial intelligence (AI), the advent of large language models (LLMs) such as GPT-3.5 Ouyang et al. (2022), GPT-4 Achiam et al. (2023), and LLaMA Touvron et al. (2023) has emerged as groundbreaking technologies, revolutionizing the field of natural language processing (NLP). LLMs have pushed the boundaries of what was previously thought possible with a “Prompt” style Liu et al. (2023a). Their capability to understand and generate human-like text has achieved state-of-the-art performance in various downstream tasks such as machine translation Manakhimova et al. (2023); Hendy et al. (2023); Khatri et al. (2023), text summarization Luo et al. (2023a); Wang et al. (2023a), and sentiment analysis Sudirjo et al. (2023); Fatouros et al. (2023). Besides, the ability of LLMs to adapt and learn from vast amounts of data has made them indispensable tools for researchers, developers, and businesses across diverse industries Fan et al. (2023); Goyal et al. (2024). Importantly, as AI continues to evolve, the impact of LLMs on our society and technology is poised to grow even further, opening up new opportunities and challenges in the realm of natural language understanding and generation.
@@
The existing LLMs are typically based on transformer architectures Vaswani et al. (2017) and trained on massive data that consists of a mixture of different languages. As shown in Figure 1, the training process of LLMs mainly consists of three stages: unsupervised pre-training, supervised fine-tuning with instructions, and aligning algorithm via reinforcement learning from human feedback (RLHF) Ouyang et al. (2022). Although LLMs have achieved remarkable advancements, their application in multilingual scenarios is still limited, especially in extremely low-resource languages Lai et al. (2023). The reason is that the language distribution of the training data for LLMs is highly imbalanced and the quality varies across languages Ding et al. (2024). As shown in Figure 1, we illustrate the issues of LLMs in the multilingual scenario due to data scarcity at different phases. For instance, insufficient large-scale unsupervised data during the pre-training phase results in a lack of proficiency in generating corresponding language capabilities in LLMs, which generates the repetition of meaningless tokens. Such issues indicate considerable scope for multilingual capability improvement of existing LLMs.
@@
To alleviate the issue of multilingualism, a common practice is to enhance the multilingual capabilities of the LLMs by incorporating corresponding multilingual data at each training stage Yang et al. (2023a); Le Scao et al. (2022); Wang et al. (2023b). However, the LLMs in the literature are mostly large-scale data-driven and face the following language issues in the multilingual scenario: (1) Knowledge Transfer: Several studies Zhang et al. (2023a); Wang et al. (2023c); Zhang et al. (2024a) have demonstrated the necessity of employing appropriate data to unleash the multilingual potential of LLMs. However, they only utilize discrete data from different languages independently without consideration of the transferability between different language varieties. Thus, the existing LLMs cannot perform well on low-resource languages Chang et al. (2023); Li et al. (2024a) and limit the number of supported languages. (2) Knowledge Accumulation: The “data island”***The data island means that data has non-existent or limited external connectivity in different storages. https://en.wikipedia.org/wiki/Data_island exists in practical scenarios due to the limitation of resource availability. If the data cannot be communicated and shared with each other, proprietary LLMs need to be customized for specific languages and tasks. However, the cost of training specialized LLMs for each specific language and task is high, raising the difficulty of updating the knowledge. Besides, the general knowledge inherited in LLMs might also be forgotten during continual training, leading to catastrophic forgetting Kirkpatrick et al. (2017). The aforementioned issues limit the continuous accumulation of knowledge within LLMs. (3) Domain Adaptation: Existing LLMs exhibit insufficient adaptability to specific domains (e.g., medical and finance) in multilingual scenarios. The domain-customized models like BioGPT Luo et al. (2022) and FinBERT Liu et al. (2021) based on domain-specific corpora are mostly English-centric. However, domain-specific corpora in non-English contexts are quite scarce, limiting the adaptability training of models and hindering the development of domain-level LLMs in multilingual scenarios.
@@
As LLMs are deeply integrated into various applications that need context comprehension and generation in multilingual scenarios, the requirement for LLMs capable of effectively and efficiently understanding and generating sequences in a language-agnostic manner becomes increasingly indispensable and urgent. Consequently, researchers have devoted significant efforts to enhancing the practicality of LLMs in multilingual scenarios from various perspectives, including training procedure (Section 3), optimization on the inference (Section 4), security in multilingual situations (Section 5) and multi-domain (Section 6). Furthermore, due to the paradigm shift from “Pre-train, Fine-tune” to “Pre-train, Prompt, Predict” in multilingual scenarios (Section 2), there has been a subtle change in the definition of multilingualism, leading to challenges in the development of the multilingual community. A series of multilingual approaches to facilitate LLMs with varying focuses have merged while lacking systematic categorization and comparison, raising the challenge of developing practical applications for specific language needs. Thus, new systematic literature and standardized definitions for introducing and comparing existing LLMs from a multilingual perspective are desirable.
@@
As the field of LLMs is developing rapidly within the AI research community, some recent surveys try to summarize these developments to provide future guidance. Xu et al. Xu et al. (2024a) propose the first survey on multilingual LLMs from three aspects, which primarily analyzes the issues of data misalignment and bias, lacking exploration from the paradigm perspective. Different from them, we conduct a structured taxonomy and comprehensive review from several perspectives. A more recent survey Qin et al. (2024) closely resembles our study, but it only classifies in terms of alignment, which is less comprehensive than ours. In particular, we not only meticulously examine the multilingual capabilities and training methods of current LLMs, but also thoroughly investigate how to uncover the potential of LLMs. In this survey, we introduce the concept of “multilingual LLMs”, and provide a comprehensive and systematic survey of existing LLMs that have remarkable multilingual capabilities. We offer categorization, comparative analysis, and multi-perspective exploration for these models, evaluating their applicability and limitations, and providing practical recommendations for their effective real-world utilization. Additionally, we discuss some useful datasets and benchmarks related to multilingualism. We also present recommendations for future research. The main contributions are as follows:
@@
Survey organization. The rest of this survey is organized as follows: In Section 2, we review the transition from pre-trained models to large language generative models in the multilingual scenario. In Section 3, we organize relevant models with various architectures, datasets, and training paradigms. Moreover, in Section 4, we investigate various multilingual inference strategies to harness the potential of LLMs for better accomplishment of multilingual tasks. In Section 5, we explore the security of LLMs with multilingual strategies, which is considered to be crucial for LLMs. In Section 6, we discuss the multi-domain issue in the multilingual real-world scenario. In addition, we present several available datasets for multilingual LLMs in Section 7 and highlight the benchmark and evaluation in Section 8. We discuss the cross-lingual bias evaluation and elimination in Section 9. Finally, we conclude this survey in Section 10.
@@ 2.1 Multilingual Models
The longstanding goal of multilingual models is to develop a universal model, capable of providing high-quality performance on any languages and tasks Huang et al. (2023a). Generally, multilingual models are trained to maximize the optimization of a mix of examples drawn from multiple language corpora Johnson et al. (2017). The benefits of multilingual models are mainly twofold. First, multilingual models can handle several languages in a single model, facilitating the knowledge transfer from high-resource languages to related but low-resource languages Huang et al. (2023b), where even languages that have never been trained before Zhang et al. (2020). Second, a single model can support multiple languages instead of training multiple language-specific models Lin et al. (2021a); Escolano et al. (2021), reducing the cost of maintaining models.
@@ 2.2 Pre-Trained Language Models
PLMs are deep neural networks initially trained on extensive unlabeled corpora, then could be adaptable to specific tasks. Existing research demonstrates that PLMs grasp and store substantial linguistic knowledge within their parameters Min et al. (2023); Li et al. (2020); Hu et al. (2023). Consequently, leveraging PLMs holds promise for enriching language comprehension and enhancing multilingual performance.
Most of the PLMs adopt the remarkable Transformer architecture Vaswani et al. (2017) as their backbone. The existing PLMs can be categorized into three typical architectures, constructed upon the Transformer encoder (e.g., BERT Devlin et al. (2018)), decoder (e.g., GPT Radford et al. (2018)), and encoder-decoder (e.g., BART Lewis et al. (2020a)), respectively. With the corresponding architecture, PLMs are usually improved based on training data sizes, parameter sizes and training strategies Li et al. (2022). For instance, GPT-3 Brown et al. (2020) is developed based on the GPT-2 Radford et al. (2019) by expanding the model size (from 1.5B to 175B) and the scale of training data, which improves performance across various language tasks. Recent research also indicates that scaling up model parameters enhances performance of PLMs Kaplan et al. (2020), leading to the emergence of large-scale PLMs such as GPT-3 (175B) Brown et al. (2020), OPT (175B) Zhang et al. (2022a), PaLM (540B) Chowdhery et al. (2023) and Switch-Transformers (1.6T) Fedus et al. (2022), boasting billions or even trillions of parameters. Besides, a new training strategy “translation language model (TLM)” is designed specifically for enhancing multilingual capabilities tasks compared to the masked language model (e.g., BERT Devlin et al. (2018)), involving translation between multiple languages Conneau and Lample (2019).

### v2

@@ Abstract
The rapid development of Large Language Models (LLMs) demonstrates remarkable multilingual capabilities in natural language processing, attracting global attention in both academia and industry. To mitigate potential discrimination and enhance the overall usability and accessibility for diverse language user groups, it is important for the development of language-fair technology. Despite the breakthroughs of LLMs, the investigation into the multilingual scenario remains insufficient, where a comprehensive survey to summarize recent approaches, developments, limitations, and potential solutions is desirable. To this end, we provide a survey with multiple perspectives on the utilization of LLMs in the multilingual scenario. We first rethink the transitions between previous and current research on pre-trained language models. Then we introduce several perspectives on the multilingualism of LLMs, including training and inference methods, information retrieval, model security, multi-domain with language culture, and usage of datasets. We also discuss the major challenges that arise in these aspects, along with possible solutions. Besides, we highlight future research directions that aim at further enhancing LLMs with multilingualism. The survey aims to help the research community address multilingual problems and provide a comprehensive understanding of the core concepts, key techniques, and latest developments in multilingual natural language processing based on LLMs.
@@ 1 Introduction
With the rapid development of artificial intelligence (AI), the advent of large language models (LLMs) such as GPT-3.5 [1], GPT-4 [2], and LLaMA [3] has emerged as groundbreaking technologies, revolutionizing the field of natural language processing (NLP). LLMs have pushed the boundaries of what was previously thought possible with a “Prompt” style [4]. Their capability to understand and generate human-like text has achieved state-of-the-art performance in various downstream tasks such as machine translation [5, 6, 7], text summarization [8, 9], and sentiment analysis [10, 11]. Besides, the ability of LLMs to adapt and learn from vast amounts of data has made them indispensable tools for researchers, developers, and businesses across diverse industries [12, 13]. Importantly, as AI continues to evolve, the impact of LLMs on our society and technology is poised to grow even further, opening up new opportunities and challenges in the realm of natural language understanding and generation.
@@
The existing LLMs are typically based on transformer architectures [14] and trained on massive data that consists of a mixture of different languages. As shown in Figure 1, the training process of LLMs mainly consists of three stages: unsupervised pre-training, supervised fine-tuning with instructions, and aligning algorithm via reinforcement learning from human feedback (RLHF) [1]. Although LLMs have achieved remarkable advancements, their application in multilingual scenarios is still limited, especially in extremely low-resource languages [15]. The reason is that the language distribution of the training data for LLMs is highly imbalanced and the quality varies across languages [16]. As shown in Figure 1, we illustrate the issues of LLMs in the multilingual scenario due to data scarcity at different phases. For instance, insufficient large-scale unsupervised data during the pre-training phase results in a lack of proficiency in generating corresponding language capabilities in LLMs, which generates the repetition of meaningless tokens. Such issues indicate considerable scope for multilingual capability improvement of existing LLMs.
@@
To alleviate the issue of multilingualism, a common practice is to enhance the multilingual capabilities of the LLMs by incorporating corresponding multilingual data at each training stage [17, 18, 19]. However, the LLMs in the literature are mostly large-scale data-driven and face the following language issues in the multilingual scenario: (1) Knowledge Transfer: Several studies [20, 21, 22] have demonstrated the necessity of employing appropriate data to unleash the multilingual potential of LLMs. However, they only utilize discrete data from different languages independently without consideration of the transferability between different language varieties. Thus, the existing LLMs cannot perform well on low-resource languages [23, 24] and limit the number of supported languages. (2) Knowledge Accumulation: The “data island”***The data island means that data has non-existent or limited external connectivity in different storages. https://en.wikipedia.org/wiki/Data_island exists in practical scenarios due to the limitation of resource availability. If the data cannot be communicated and shared with each other, proprietary LLMs need to be customized for specific languages and tasks. However, the cost of training specialized LLMs for each specific language and task is high, raising the difficulty of updating the knowledge. Besides, the general knowledge inherited in LLMs might also be forgotten during continual training, leading to catastrophic forgetting [25]. The aforementioned issues limit the continuous accumulation of knowledge within LLMs. (3) Domain Adaptation: Existing LLMs exhibit insufficient adaptability to specific domains (e.g., medical and finance) in multilingual scenarios. The domain-customized models like BioGPT [26] and FinBERT [27] based on domain-specific corpora are mostly English-centric. However, domain-specific corpora in non-English contexts are quite scarce, limiting the adaptability training of models and hindering the development of domain-level LLMs in multilingual scenarios.
@@
As LLMs are deeply integrated into various applications that need context comprehension and generation in multilingual scenarios, the requirement for LLMs capable of effectively and efficiently understanding and generating sequences in a language-agnostic manner becomes increasingly indispensable and urgent. Consequently, researchers have devoted significant efforts to enhancing the practicality of LLMs in multilingual scenarios from various perspectives, including training procedure (Section 3), optimization on the inference (Section 4), information retrieval systems (Section 5), security in multilingual situations (Section 6) and multi-domain (Section 7). Furthermore, due to the paradigm shift from “Pre-train, Fine-tune” to “Pre-train, Prompt, Predict” in multilingual scenarios (Section 2), there has been a subtle change in the definition of multilingualism, leading to challenges in the development of the multilingual community. A series of multilingual approaches to facilitate LLMs with varying focuses have merged while lacking systematic categorization and comparison, raising the challenge of developing practical applications for specific language needs. Thus, new systematic literature and standardized definitions for introducing and comparing existing LLMs from a multilingual perspective are desirable.
@@
As the field of LLMs is developing rapidly within the AI research community, some recent surveys try to summarize these developments to provide future guidance. Xu et al. [28] propose the first survey on multilingual LLMs from three aspects, which primarily analyzes the issues of data misalignment and bias, lacking exploration from the paradigm perspective. Different from them, we conduct a structured taxonomy and comprehensive review from several perspectives. A more recent survey [29] closely resembles our study, but it only classifies in terms of alignment, which is less comprehensive than ours. In particular, we not only meticulously examine the multilingual capabilities and training methods of current LLMs, but also thoroughly investigate how to uncover the potential of LLMs. In this survey, we introduce the concept of “multilingual LLMs”, and provide a comprehensive and systematic survey of existing LLMs that have remarkable multilingual capabilities. We offer categorization, comparative analysis, and multi-perspective exploration for these models, evaluating their applicability and limitations, and providing practical recommendations for their effective real-world utilization. Additionally, we discuss some useful datasets and benchmarks related to multilingualism. We also present recommendations for future research. The main contributions are as follows:
@@
Survey organization. The rest of this survey is organized as follows: In Section 2, we review the transition from pre-trained models to large language generative models in the multilingual scenario. In Section 3, we organize relevant models with various architectures, datasets, and training paradigms. Moreover, in Section 4, we investigate various multilingual inference strategies to harness the potential of LLMs for better accomplishment of multilingual tasks. In Section 5, we provide a preliminary exploration of the integration of multilingual information retrieval systems and LLMs, aiming to reveal the development opportunities for multilingual information retrieval systems in the era of LLMs. In Section 6, we explore the security of LLMs with multilingual strategies, which is considered to be crucial for LLMs. In Section 7, we discuss the multi-domain issue in the multilingual real-world scenario. In addition, we present several available datasets for multilingual LLMs in Section 8 and highlight the benchmark and evaluation in Section 9. We discuss the cross-lingual bias evaluation and elimination in Section 10. Finally, we conclude this survey in Section 11.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, calign=edge midpoint, , [LLMs with Multilingualism, text width=1.7cm, fill=black!10 [Representative Models §3, text width=1.8cm, for tree=fill=red!20 [Training from scratch §3.1, text width=1.8cm [ GPT-3 [30]; mT5 [31]; ByT5 [32]; Gopher [33]; LaMDA [34]; OPT [35]; PaLM [36]; mGPT [37]; GPT-3.5 [1]; XGLM [38]; LLaMA [3]; GPT-4 [2]; PANGU-$\sum$ [39]; Pythia [40]; PaLM-2 [41]; InternLM [42]; PolyLM [43]; LLaMA-2 [3]; Baichuan-2 [44]; Qwen [45]; Mistral [46]; Gemini [47]; TigerBot [48]; YAYI-2 [49]; DeepSeek [50]; Orion [51]; TeleChat [52]; Claude3 [53]; InternLM2 [54]; Phi-3 [52]; LLaMA-3 [55], text width=7.0cm, node options=align=left ] ] [Continual Training §3.2, text width=1.8cm [ FLAN-T5 [56]; FLAN-PaLM [56]; ChatGLM [57]; Alpaca [58]; ParroT [59]; BigTrans [17]; Vicuna [60]; BayLing [20]; OpenChat [21]; GLM-4 [61]; Aya [62], text width=7.0cm, node options=align=left ] ] ] [Inference Strategies §4, text width=1.8cm, for tree=fill=yellow!20 [Translation-Based §4.2, text width=1.8cm [ Intrator et al. [63]; Liu et al. [64], text width=7.0cm, node options=align=left ] ] [Chain-of-Thought §4.3, text width=1.8cm [ Liu et al. [64]; XLT [65]; Shi et al. [66]; Kim et al. [67]; Suzgun et al. [68]; xCoT [69], text width=7.0cm, node options=align=left ] ] [Code-Switching §4.4, text width=1.8cm [ Zhang et al. [70]; Koto et al. [71]; Peng et al. [72], text width=7.0cm, node options=align=left ] ] [Retrieval Augmented Generation §4.5, text width=1.8cm [ Zhang et al. [73]; Shi et al. [74]; Agrawal et al. [75]; Li et al. [76]; Li et al. [77]; Winata et al. [78]; Garcia et al. [79]; Ramos et al. [80]; Kim et al. [81]; Thakur et al. [82]; Sennrich et al. [83]; Vernikos et al. [84]; Fu et al. [85]; Zeng et al. [86]; He et al. [87]; Conia et al. [88], text width=7.0cm, node options=align=left ] ] ] [Information Retrieval §5, text width=1.8cm, for tree=fill=orange!20 [Training Data §5.1, text width=1.8cm [ InPars [89]; InPars-v2 [90]; InPars Toolkit [91]; Promptagator [92]; SwimIR [93]; JH-POLO [94]; mE5-Mistral [95]; Gecko [96]; Arctic-Embed [97], text width=7.0cm, node options=align=left ] ] [Retrievers §5.2, text width=1.8cm [ mE5 [98]; mGTE [99]; BGE [100]; OpenAI-Embed [101]; Cohere-Embed [102]; voyage-multilingual-2 [103]; RepLLAMA [104]; PromptReps [105]; MTEB [106]; NV-Embed [107]; Springer et al. [108]; LLM2Vec [109]; Muennighoff [110]; Kusupati et al. [111]; LM-Cocktail [112]; AnglE [113]; NLLB-E5 [114]; mColBERT [115, 116]; ColBERT-X [117]; Translate-Distill [118]; Huang et al. [119]; Lawrie et al. [120]; Yang et al. [121]; ColBERT-XM [122]; BGE-M3 [123], text width=7.0cm, node options=align=left ] ] [Rerankers §5.3, text width=1.8cm [ mMARCO [115]; Jeronymo et al. [124] RankGPT [125]; LRL [126]; Setwise [127]; RankVicuna [128]; RankZephyr [129]; Rank-without-GPT [130]; TourRank [131]; Adeyemi et al. [132], text width=7.0cm, node options=align=left ] ] ] [Security §6, text width=1.8cm, for tree=fill=blue!20 [Attack (Jailbreak) §6.1, text width=1.0cm, for tree=fill=blue!20 [Greedy Coordinate Gradient, text width=1.4cm [ Sitawarin et al. [133]; Zou et al. [134], text width=6.0cm, node options=align=left ] ] [Prompt-Based, text width=1.4cm [ Wei et al. [135]; Liu et al. [136]; Shen et al. [137]; Deng et al. [138]; Li et al. [139]; Liu et al. [140]; Jin et al. [141], text width=6.0cm, node options=align=left ] ] [Multilingual, text width=1.4cm [ Shen et al. [142]; Deng et al. [143]; Puttaparthi et al. [144]; Yong et al. [145]; Xu et al. [146]; Li et al. [147]; Yuan et al. [148]; Huang et al. [149], text width=6.0cm, node options=align=left ] ] ] [Defense §6.2, text width=1.0cm, for tree=fill=blue!20 [Open-Source Models, text width=1.4cm [ Robey et al. [150]; Deng et al. [143]; Li et al. [147]; Zhou et al. [151]; GUARD [141], text width=6.0cm, node options=align=left ] ] [Closed-Source Models, text width=1.4cm [ Jain et al. [152]; Wu et al. [153]; Li et al. [154], text width=6.0cm, node options=align=left ] ] ] ] [Multidomain §7, text width=1.8cm, for tree=fill=green!20 [Medical §7.1, text width=1.8cm [ KBioXLM [155]; BioMistral [156]; MMedLM2 [157]; Apollo [158]; L2M3 [159]; Medical mT5 [160], text width=7.0cm, node options=align=left ] ] [Legal §7.2, text width=1.8cm [ LEXTREME [161]; Brugger et al. [162]; Christen et al. [163]; Baumgartner et al. [164]; Niklaus et al. [165]; LegalLAMA [166]; Trautmann et al. [167], text width=7.0cm, node options=align=left ] ] ] [Datasets, text width=1.8cm, for tree=fill=lightblue!20 [Corpus §8, text width=1.8cm [ Amazon intent [168]; Amazon reviews [169] ; Aya [170]; Bactrian-x [171]; Biblenlp [172]; Bloom-lm [18]; CC100 [173] [174]; CulturaX [175]; GPT-4 Prompts [176] ; Guanaco [177]; HPLT [178] ; IWSLT 2017 [179]; mC4 [180]; Mewsli-x [181]; Minds14 [182] ; Miracl [183] ; MLDR [184]; MMedC [185]; MQA [186]; Multi-sentiments [187]; Multiconer [188] [189]; Nomiracl [190] [82];Open Subtitles [191] ; OSCAR [192]; Para-pat [193] ; Project Gutenberg [194]; ShareGPT [195]; SREDFM [196]; TED Talks [197]; TED-talks-iwslt [198]; Toxi-text [199]; UD [200]; Wikiann [201] [202]; Wikipedia [203]; Wit Base [204]; xP3 [205], text width=7.0cm, node options=align=left ] ] [Benchmark §9, text width=1.8cm [ Afrisent [206]; ASPEN [207] ; BELEBELE [208]; BioInstructQA [209]; Bucc-bitext-mining [210]; Cross-Sum [211]; Crossmodal-3600 [212]; Exams [213]; Fairlex [214]; FLORES-200 [215] [216] [217]; GEOMLAMA [218]; Humaneval-XL [219]; M-Hellaswag [220]; M-MMLU [220]; M3Exam [221]; M3LS [222]; MARC [223]; MasakhaNER [224]; Masakhanews [225]; MASSIVE [226]; MaXM [227]; MEGA [228]; MEGAVerse [229]; Mela [230]; MGSM [66]; MLQA [231]; MMedBench [232]; Multi-CoNER [233]; MULTIEURLEXDOC [234]; Multilingual-Fig-QA [235]; NusaX [236]; ODEX [237]; OPUS-100 [238]; Paws-X [239]; Pmindiasum [240]; PRESTO [241]; SEAHORSE [242]; Sib200 [243]; SMiLER [244]; STSB-multi-mt [245]; Tatoeba-mt [246]; Tydip [247]; TyDiQA [248]; Universal Dependencies [200] ; X-CLAIM [249]; X-RiSAWOZ [250]; XCOPA [251]; XCSQA [252]; xDial-Eval [253]; XGLUE [254]; XL-SUM [255]; XNLI [256]; XQuAD [257]; XSEMPLR [258]; XStoryCloze [38]; XTREME [259]; XTREME-R [260]; XWinograd [261]; XCSR [262], text width=7.0cm, node options=align=left ] ] ] ]
@@ 2.1 Multilingual Models
The longstanding goal of multilingual models is to develop a universal model, capable of providing high-quality performance on any languages and tasks [263]. Generally, multilingual models are trained to maximize the optimization of a mix of examples drawn from multiple language corpora [264]. The benefits of multilingual models are mainly twofold. First, multilingual models can handle several languages in a single model, facilitating the knowledge transfer from high-resource languages to related but low-resource languages [265], where even languages that have never been trained before [238]. Second, a single model can support multiple languages instead of training multiple language-specific models [266, 267], reducing the cost of maintaining models.
@@ 2.2 Pre-Trained Language Models
PLMs are deep neural networks initially trained on extensive unlabeled corpora, then could be adaptable to specific tasks. Existing research demonstrates that PLMs grasp and store substantial linguistic knowledge within their parameters [268, 269, 270]. Consequently, leveraging PLMs holds promise for enriching language comprehension and enhancing multilingual performance.
Most of the PLMs adopt the remarkable Transformer architecture [14] as their backbone. The existing PLMs can be categorized into three typical architectures, constructed upon the Transformer encoder (e.g., BERT [271]), decoder (e.g., GPT [272]), and encoder-decoder (e.g., BART [273]), respectively. With the corresponding architecture, PLMs are usually improved based on training data sizes, parameter sizes and training strategies [274]. For instance, GPT-3 [30] is developed based on the GPT-2 [275] by expanding the model size (from 1.5B to 175B) and the scale of training data, which improves performance across various language tasks. Recent research also indicates that scaling up model parameters enhances performance of PLMs [276], leading to the emergence of large-scale PLMs such as GPT-3 (175B) [30], OPT (175B) [35], PaLM (540B) [36] and Switch-Transformers (1.6T) [277], boasting billions or even trillions of parameters. Besides, a new training strategy “translation language model (TLM)” is designed specifically for enhancing multilingual capabilities tasks compared to the masked language model (e.g., BERT [271]), involving translation between multiple languages [278].